OpenStudy (anonymous):

Need HELP ASAP, about 6 problems i dont understand

OpenStudy (anonymous):

Question 7 (Multiple Choice Worth 3 points) Complete the Next Step: A partial solution to the equation 4d − 12 = 12 − 4d is given below. Choose the step that is missing from the solution. 4d − 12 = 12 − 4d 4d − 24 = −4d 3 = d 24 = −8d −24 = 8d −24 = 0d −24 = −8d Question 8 (Multiple Choice Worth 3 points) Solve: −3x − 5 = 16 −3.7 3 −7 7 Question 9 (Multiple Choice Worth 3 points) Solve: 2(x + 1) = −8 5 −3 −5 −4.5 Question 10 (Multiple Choice Worth 3 points) 3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property Question 11 (Multiple Choice Worth 3 points) Solve: x + 1 = −x + 8 7 14 18 Question 12 (Multiple Choice Worth 3 points) Solve: 2(x + 2) − (x + 1) = 8 2 3 4 5 Question 13 (Multiple Choice Worth 3 points) Solve: 6 − x = 15 −21 21 −9 9 Question 14 (Essay Worth 6 points) Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer. Question 15 (Essay Worth 6 points) What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

Do you know how to solve equations? You always do the same thing to both sides, e.g. add 5 to both sides.

OpenStudy (anonymous):

so if it says x-5=4 then you add 5 to both sides of the "=" to get x=9

OpenStudy (anonymous):

yeah i understand that but its complicated to me, i have not done any math for a while now

OpenStudy (anonymous):

ok so in the first question (Q7) the first step was to take away 12 from both sides. Then the next step would be to take away 4d from both sides so all the d's are on the same side of the "=".

OpenStudy (anonymous):

Okay so it will be 3?

OpenStudy (anonymous):

no because -4d-4d is not 0d. what is -4-4?

OpenStudy (anonymous):

-8